# CLIP4STR: A Simple Baseline for Scene Text Recognition with Pre-trained Vision-Language Model

### v3

@@ Abstract
Pre-trained vision-language models (VLMs) are the de-facto foundation models for various downstream tasks. However, scene text recognition methods still prefer backbones pre-trained on a single modality, namely, the visual modality, despite the potential of VLMs to serve as powerful scene text readers. For example, CLIP can robustly identify regular (horizontal) and irregular (rotated, curved, blurred, or occluded) text in images. With such merits, we transform CLIP into a scene text reader and introduce CLIP4STR, a simple yet effective STR method built upon image and text encoders of CLIP. It has two encoder-decoder branches: a visual branch and a cross-modal branch. The visual branch provides an initial prediction based on the visual feature, and the cross-modal branch refines this prediction by addressing the discrepancy between the visual feature and text semantics. To fully leverage the capabilities of both branches, we design a dual predict-and-refine decoding scheme for inference. We scale CLIP4STR in terms of the model size, pre-training data, and training data, achieving state-of-the-art performance on 11 STR benchmarks. Additionally, a comprehensive empirical study is provided to enhance the understanding of the adaptation of CLIP to STR. We believe our method establishes a simple yet strong baseline for future STR research with VLMs.
@@ I Introduction
Vision-language models (VLMs) pre-trained on web-scale data like CLIP [1] and ALIGN [2] shows remarkable zero-shot capacity across different tasks. Researchers also successfully transfer the knowledge from pre-trained VLMs to diverse tasks in a zero-shot or fine-tuning manner, e.g., visual question answering [3], image manipulation [4, 5], information retrieval [6, 7], referring expression comprehension [8], and image captioning [9]. Currently, VLM is widely recognized as a foundational model and an important component of artificial intelligence [10].
@@
STR methods generally struggle with irregular text like rotated, curved, blurred, or occluded text [15, 16]. However, irregular text is prevalent in real-life scenarios [17, 18], making it necessary for STR models to effectively handle these challenging cases. Interestingly, we observe that the VLM (e.g., CLIP [1]) can robustly perceive irregular text in natural images. In Figure 1, we put different text stickers on a natural image, use CLIP to classify it [^1] The class categories are from CIFAR-10 [19]. The experiment is inspired by Stanislav Fort [20]. , and visualize the attention of CLIP via Grad-CAM [21]. It is evident that CLIP pays high attention to the text sticker and accurately understands the meaning of the word, regardless of text variations [^2] This phenomenon, where CLIP focuses on the text while disregarding the natural object, is also known as typographic attacks [22]. Neurons in CLIP image encoders can simultaneously perceive both visual and text signals associated with the same concept, such as an image or typographic text of Spiderman. This ability may stem from the training images containing scene texts in the large training data. . CLIP is trained on massive natural images collected from the web and its text perception ability may come from the natural images containing scene texts [22]. Will CLIP perceive the text in common STR images [23, 24, 18], which are cropped from a natural image? Figure 2 presents the visualization results of CLIP-ViT-B/32 for STR images. Although the text in these STR images is occluded, curved, blurred, and rotated, CLIP can still perceive them. From Figure 1&2, we can see CLIP possesses an exceptional capability to perceive and comprehend various text in images. This is exactly the desired quality for a robust STR backbone.
@@
We scale CLIP4STR across different model sizes, pre-training data, and training data to investigate the effectiveness of large scale pre-trained VLMs as STR backbones. CLIP4STR achieves state-of-the-art performance on 11 STR benchmarks, encompassing both regular and irregular text. Additionally, we present a comprehensive empirical study on adapting CLIP to STR. We believe CLIP4STR provides a simple but strong baseline for future STR research with VLMs.
@@ II-A Vision-Language Models and Its Application
Large-scale pre-trained vision-language models learning under language supervision such as CLIP [1], ALIGN [2], and Florence [27] demonstrate excellent generalization abilities. This encourages researchers to transfer the knowledge of these pre-trained VLMs to different downstream tasks in a fine-tuning or zero-shot fashion. For instance, [6, 28, 29] tune CLIP on videos and make CLIP specialized in text-video retrieval, DiffusionCLIP [5] introduces CLIP to zero-shot image manipulation, CLIPScore [9] uses CLIP to evaluate the quality of generated image captions, and [30, 31] use CLIP as the reward model during test time or training. The wide application of VLMs also facilitates the research on different pre-training models, e.g., ERNIE-ViLG [32], CoCa [33], OFA [34], DeCLIP [35], FILIP [36], and ALBEF [37]. Researchers also try to explore the power of scaling up the data, e.g., COYO-700M [38] and LAION-5B [39]. Generally, more data brings more power for large VLMs [40].
VLMs pre-trained on large-scale image-text pairs possess many fascinating attributes [1, 22, 41]. For instance, some neurons in CLIP can perceive the visual and text signals corresponding to the same concept. [22] finds particular neurons in CLIP-RN50$\times$4 respond to both photos of Spiderman and the text "spider" in an image. This also leads to Typographic Attacks, namely, VLMs focus on the text rather than natural objects in an image as shown in Figure 1. In this work, we leverage the text perception ability of multi-modal neurons and make CLIP specialize in scene text recognition.
@@ II-B Scene Text Recognition
Scene text recognition methods can be broadly divided into two categories: context-free and context-aware. Context-free STR methods only utilize the visual features of images, such as CTC-based [42] methods [43, 44, 45, 12], segmentation-based methods [46, 47, 48], and attention-based methods with an encoder-decoder mechanism [49, 50]. Since context-free STR methods lack the understanding of text semantics, they are less robust against occluded or incomplete text. Context-aware STR methods are the mainstream approach now, leveraging text semantics to enhance recognition performance. For example, ABINet [11], LevOCR [51], MATRN [52], and TrOCR [14] incorporate an external language model to capture text semantics. Other methods achieve similar goals with built-in modules, such as RNN [53, 54], GRU [55], transformer [56, 26, 57]. The context information is interpreted as the relations of textual primitives by Zhang et al[58], who proposes a relational contrastive self-supervised learning STR framework. Besides the context-free and context-aware methods, some efforts aim to enhance the explainability of STR. For instance, STRExp [59] utilizes local individual character explanations to deepen the understanding of STR methods. Moreover, training data plays a vital role in STR. Traditionally, synthetic data [60, 61] are used for training due to the ease of generating a large number of samples. However, recent research suggests that using realistic training data can lead to better outcomes compared to synthetic data [62, 26, 63, 64]. Motivated by these findings, we primarily employ realistic training data in this work.
The success of VLMs also spreads to the STR area. For example, TrOCR [14] adopts separate pre-trained language and vision models plus post-pretraining on STR data in an auto-regressive manner [65], MATRN [52] uses a popular multi-modal fusion manner in VLMs such as ALBEF [37] and ViLT [66]. CLIPTER [67] enhances the character recognition performance by utilizing the CLIP features extracted from the global image. CLIP-OCR [68] leverages both visual and linguistic knowledge from CLIP through feature distillation. In contrast, we directly transfers CLIP to a robust scene text reader, eliminating the need for CLIP features from the global image or employing an additional CLIP model as a teacher for the STR reader. We hope that our method can serve as a strong baseline for future STR research with VLMs.
@@ III-A1 CLIP
CLIP consists of a text encoder and an image encoder. CLIP is pre-trained on 400 million image-text pairs using contrastive learning. The text and image features from CLIP are aligned in a joint image-text embedding space. i) The image encoder of CLIP is a vision transformer (ViT) [69]. Given an image, ViT introduces a visual tokenizer (convolution) to convert non-overlapped image patches into a discrete sequence. A [CLASS] token is then prepended to the beginning of the image sequence. Initially, CLIP image encoder only returns the feature of the [CLASS] token, but in this work, we return features of all tokens. The rationale behind this choice is that character-level recognition requires fine-grained detail, and local features from all patches are necessary. These features are normalized and linearly projected into the joint image-text embedding space. ii) The text encoder of CLIP is a transformer encoder [25, 70]. The text tokenizer is a lower-cased byte pair encoding – BPE [71] with vocabulary size 49,152. The beginning and end of the text sequence are padded with [SOS] and [EOS] tokens, respectively. Linguistic features of all tokens are ultilized for character recoginiton. These features are also normalized and linearly projected into the joint image-text embedding space.
@@ III-A2 Permuted sequence modeling
Traditionally, STR methods use a left-to-right or right-to-left order to model character sequences  [11]. However, the characters in a word do not strictly follow such directional dependencies. For instance, to predict the letter ”o” in the word ”model”, it is sufficient to consider only the context ”m_de” rather than relying solely on the left-to-right context ”m_” or the right-to-left context ”led_”. The dependencies between characters in a word can take various forms. To encourage the STR method to explore these structural relationships within words, PARSeq [26] introduces a permuted sequence modeling (PSM) technique. This technique uses a random attention mask $\mathcal{M}$ for attention operations [25] to generate random dependency relationships between the input context and the output. Table I illustrates three examples of mask $\mathcal{M}$. We will delve further into this mechanism in Section III-C.
Figure 4: The decoder of CLIP4STR. [B], [E], and [P] are the beginning, end, and padding tokens, respectively. []delimited-[][$\cdots$][ ⋯ ]’ in prediction represents the ignored outputs. Layer normalization [72] and dropout [73] in the decoder are ignored.
@@ III-B Encoder
The framework of CLIP4STR is illustrated in Figure 3. CLIP4STR employs a dual encoder-decoder design, consisting of a visual branch and a cross-modal branch. The text and image encoders utilize the architectures and pre-trained weights from CLIP. The visual branch generates an initial prediction based on the visual features extracted by the image encoder. Subsequently, the cross-modal branch refines the initial prediction by addressing the discrepancy between the visual features and the textual semantics of the prediction. Since the image and text features are aligned in a joint image-text embedding space during pre-training, it becomes easy to identify this discrepancy. The cross-modal branch acts as a semantic-aware spell checker.
@@ III-C Decoder
The decoder aims to extract the character information from the visual feature $\bm{F}_{i}$ or cross-modal feature $\bm{F}_{c}$. The decoder framework is shown in Figure 4. It adopts the design of the transformer decoder [25] plus the PSM technique mentioned in Section III-A2, enabling a predicted character to have arbitrary dependencies on the input context during training.
The visual and cross-modal decoders have the same architecture but differ in the input. They receive the following inputs: a learnable position query $\bm{p}\in\mathbb{R}^{N\times D}$, an input context $\bm{c}\in\mathbb{R}^{N\times D}$, and a randomly generated attention mask $\mathcal{M}\in\mathbb{R}^{N\times N}$. $N$ represents the length of characters. The decoder outputs the prediction $\bm{y}\in\mathbb{R}^{N\times C}$, where $C$ is the number of character classes. The decoding stage can be denoted as
$\bm{y}=\texttt{DEC}(\bm{p},\bm{c},\mathcal{M},\bm{F}).$ (4)
The first Multi-Head Attention (MHA) in Figure 4 performs context-position attention:
@@
During training, the output of the decoder depends on the input context in an arbitrary manner. This encourages the decoder to analyze the word structure beyond the traditional left-to-right or right-to-left sequence modeling assumptions [11]. The inclusion of a random attention mask $\mathcal{M}$ in Eq.(5) enables this capability [26]. Table I presents examples of generated attention masks, including a left-to-right auto-regressive (AR) mask, a cloze mask, and a random mask. Following PARSeq [26], we employ $K=6$ masks per input context during training. The first two masks are left-to-right and right-to-left masks, and others are randomly generated.
CLIP4STR is optimized to minimize the sum of cross-entropy losses ($\texttt{CE}(\cdot)$) of the visual branch and the cross-modal branch:
@@
where $\hat{\bm{y}}$ represents the ground truth, $\bm{y}^{i}$ is the prediction of the visual branch, and $\bm{y}$ is the prediction of the cross-modal branch.
@@ 22       $\bm{y}\leftarrow\texttt{Dec}^{c}(\bm{p}^{c},\bm{c},\mathcal{M}^{c},\bm{F}_{c})$;
Algorithm 1 Inference decoding scheme, detailed explanation refer to Appendix B
@@ III-C1 Decoding scheme
CLIP4STR consists of two branches: a visual branch and a cross-modal branch. To fully exploit the capacity of both branches, we design a dual predict-and-refine decoding scheme for inference, inspired by previous STR methods [11, 26]. Algorithm 1 illustrates the decoding process. The visual branch first performs autoregressive decoding, where the future output depends on previous predictions. Subsequently, the cross-modal branch addresses possible discrepancies between the visual feature and the text semantics of the visual prediction, aiming to improve recognition accuracy. This process is also autoregressive. Finally, the previous predictions are utilized as the input context for refining the output in a cloze-filling manner. The refinement process can be iterative. After iterative refinement, the output of the cross-modal branch serves as the final prediction.
@@
TABLE III: Word accuracy on 8 common benchmarks. Thebestandsecond-bestresults are highlighted. Benchmark datasets (B) - SVT, IIIT5K, IC13, and IC15. ‘N/A’ for not applicable. Reproduced by PARSeq [26].
Method Pre-train Data Train Data IIIT5K SVT IC13 IC15 IC15 SVTP CUTE HOST WOST
3,000 647 1,015 1,811 2,077 645 288 2,416 2,416

### v4

@@ Abstract
Pre-trained vision-language models (VLMs) are the de-facto foundation models for various downstream tasks. However, scene text recognition methods still prefer backbones pre-trained on a single modality, namely, the visual modality, despite the potential of VLMs to serve as powerful scene text readers. For example, CLIP can robustly identify regular (horizontal) and irregular (rotated, curved, blurred, or occluded) text in images. With such merits, we transform CLIP into a scene text reader and introduce CLIP4STR, a simple yet effective STR method built upon image and text encoders of CLIP. It has two encoder-decoder branches: a visual branch and a cross-modal branch. The visual branch provides an initial prediction based on the visual feature, and the cross-modal branch refines this prediction by addressing the discrepancy between the visual feature and text semantics. To fully leverage the capabilities of both branches, we design a dual predict-and-refine decoding scheme for inference. We scale CLIP4STR in terms of the model size, pre-training data, and training data, achieving state-of-the-art performance on 13 STR benchmarks. Additionally, a comprehensive empirical study is provided to enhance the understanding of the adaptation of CLIP to STR. Our method establishes a simple yet strong baseline for future STR research with VLMs.
@@ I Introduction
Vision-language models (VLMs) pre-trained on web-scale data like CLIP [1] and ALIGN [2] shows remarkable zero-shot capacity across different tasks. Researchers also successfully transfer the knowledge from pre-trained VLMs to diverse tasks in a zero-shot or fine-tuning manner, e.g., visual question answering [3], information retrieval [4, 5], referring expression comprehension [6], and image captioning [7]. VLM is widely recognized as a foundational model and an important component of artificial intelligence [8].
@@
STR methods generally struggle with irregular text like rotated, curved, blurred, or occluded text [13, 14]. However, irregular text is prevalent in real-life scenarios [15, 16], making it necessary for STR models to effectively handle these challenging cases. Interestingly, we observe that the VLM (e.g., CLIP [1]) can robustly perceive irregular text in natural images. In Fig. 1, we put different text stickers on a natural image, use CLIP to classify it [^1] The class categories are from CIFAR-10 [17]. The experiment is inspired by Stanislav Fort [18]. , and visualize the attention of CLIP via Grad-CAM [19]. It is evident that CLIP pays high attention to the text sticker and accurately understands the meaning of the word, regardless of text variations [^2] This phenomenon, where CLIP focuses on the text while disregarding the natural object, is also known as typographic attacks [20]. Neurons in CLIP image encoders can simultaneously perceive both visual and text signals associated with the same concept, such as an image or typographic text of Spiderman. This ability may stem from the training images containing scene texts in the large training data. . CLIP is trained on massive natural images collected from the web, and its text perception ability may come from the natural images containing scene texts [20]. Will CLIP perceive the text in common STR images [21, 22, 16], which are cropped from a natural image? Fig. 2 presents the visualization results of CLIP-ViT-B/32 for STR images. Although the text in these STR images is occluded, curved, blurred, and rotated, CLIP can still perceive them. From Fig. 1&2, we can see CLIP possesses an exceptional capability to perceive and comprehend various text in images. This is exactly the desired quality for a robust STR backbone.
@@
We scale CLIP4STR across different model sizes, pre-training data, and training data to investigate the effectiveness of large-scale pre-trained VLMs as STR backbones. CLIP4STR achieves state-of-the-art performance on 13 STR benchmarks, encompassing regular and irregular text. Additionally, we present a comprehensive empirical study on adapting CLIP to STR. CLIP4STR provides a simple yet strong baseline for future STR research with VLMs.
@@ II-A Vision-Language Models and Its Application
Large-scale pre-trained vision-language models learning under language supervision such as CLIP [1], ALIGN [2], and Florence [25] demonstrate excellent generalization abilities. This encourages researchers to transfer the knowledge of these pre-trained VLMs to different downstream tasks in a fine-tuning or zero-shot fashion. For instance, [4, 26, 27] tune CLIP on videos and make CLIP specialized in text-video retrieval, CLIPScore [7] uses CLIP to evaluate the quality of generated image captions, and [28, 29] use CLIP as the reward model during test time or training. The wide application of VLMs also facilitates the research on different pre-training models, e.g., ERNIE-ViLG [30], CoCa [31], OFA [32], DeCLIP [33], FILIP [34], and ALBEF [35]. Researchers also explore the power of scaling up the data, e.g., COYO-700M [36] and LAION-5B [37]. Generally, more data brings more power for large VLMs [38].
VLMs pre-trained on large-scale image-text pairs possess many fascinating attributes [1, 20, 39]. For instance, some neurons in CLIP can perceive the visual and text signals corresponding to the same concept. [20] finds particular neurons in CLIP-RN50$\times$4 respond to both photos of Spiderman and the text ‘‘spider’’ in an image. This also leads to Typographic Attacks, namely, VLMs focus on the text rather than natural objects in an image as shown in Figure 1. In this work, we leverage the text perception ability of multi-modal neurons and make CLIP specialize in scene text recognition.
@@ II-B Scene Text Recognition
Scene text recognition methods can be broadly divided into two categories: context-free and context-aware. Context-free STR methods only utilize the visual features of images, such as CTC-based [40] methods [41, 42, 43, 10], segmentation-based methods [44, 45, 46], and attention-based methods with an encoder-decoder mechanism [47, 48]. Since context-free STR methods lack the understanding of text semantics, they are less robust against occluded or incomplete text. Context-aware STR methods are the mainstream approach, leveraging text semantics to enhance recognition performance. For example, ABINet [9], LevOCR [49], MATRN [50], and TrOCR [12] incorporate an external language model to capture text semantics. Other methods achieve similar goals with built-in modules, such as RNN [51, 52], GRU [53], transformer [54, 24, 55]. The context information is interpreted as the relations of textual primitives by Zhang et al[56], who proposes a relational contrastive self-supervised learning STR framework. Besides the context-free and context-aware methods, some efforts aim to enhance the explainability of STR. For instance, STRExp [57] utilizes local individual character explanations to deepen the understanding of STR methods. Moreover, training data plays a vital role in STR. Traditionally, synthetic data [58, 59] are used for training due to the ease of generating a large number of samples. However, recent research suggests that using realistic training data can lead to better outcomes compared to synthetic data [60, 24, 61, 62]. Motivated by these findings, we primarily employ realistic training data in this work.
The success of VLMs also spreads to the STR area. For example, TrOCR [12] adopts separate pre-trained language and vision models plus post-pretraining on STR data in an auto-regressive manner [63], MATRN [50] uses a popular multi-modal fusion manner in VLMs such as ALBEF [35] and ViLT [64]. CLIPTER [65] enhances the character recognition performance by utilizing the CLIP features extracted from the global image. CLIP-OCR [66] leverages both visual and linguistic knowledge from CLIP through feature distillation. In contrast, we directly transfer CLIP to a robust scene text reader, eliminating the need for CLIP features from the global image or employing an additional CLIP model as a teacher for the STR reader. We hope our method can be a strong baseline for future STR research with VLMs.
@@ III-A Preliminary
Figure 4: The decoder of CLIP4STR. [B], [E], and [P] are the beginning, end, and padding tokens, respectively. ‘[]delimited-[][$\cdots$][ ⋯ ]’ in prediction represents the ignored outputs. Layer normalization [67] and dropout [68] in the decoder are ignored.
@@ III-A1 CLIP
CLIP consists of a text encoder and an image encoder. CLIP is pre-trained on 400 million image-text pairs using contrastive learning. The text and image features from CLIP are aligned in a joint image-text embedding space. i) The image encoder of CLIP is a vision transformer (ViT) [69]. Given an image, ViT introduces a visual tokenizer (convolution) to convert non-overlapped image patches into a discrete sequence. A [CLASS] token is then prepended to the beginning of the image sequence. Initially, CLIP image encoder only returns the feature of the [CLASS] token, but in this work, we return features of all tokens. The rationale behind this choice is that character-level recognition requires fine-grained detail, and local features from all patches are necessary. These features are normalized and linearly projected into the joint image-text embedding space. ii) The text encoder of CLIP is a transformer encoder [23, 70]. The text tokenizer is a lower-cased byte pair encoding – BPE [71] with vocabulary size 49,152. The beginning and end of the text sequence are padded with [SOS] and [EOS] tokens, respectively. Linguistic features of all tokens are utilized for character recognition. These features are also normalized and linearly projected into the joint image-text embedding space.
@@ III-A2 Permuted sequence modeling
Traditionally, STR methods use a left-to-right or right-to-left order to model character sequences [9]. However, the characters in a word do not strictly follow such directional dependencies. For instance, to predict the letter “o” in the word “model”, it is sufficient to consider only the context “m_de” rather than relying solely on the left-to-right context “m_” or the right-to-left context “led_”. The dependencies between characters in a word can take various forms. To encourage the STR method to explore these structural relationships within words, PARSeq [24] introduces a permuted sequence modeling (PSM) technique. This technique uses a random attention mask $\mathcal{M}$ for attention operations [23] to generate random dependency relationships between the input context and the output. Table I illustrates three examples of mask $\mathcal{M}$. We will delve further into this mechanism in §III-C.
@@ III-B Encoder
The framework of CLIP4STR is illustrated in Fig. 3. CLIP4STR employs a dual encoder-decoder design, consisting of a visual branch and a cross-modal branch. The text and image encoders utilize the architectures and pre-trained weights from CLIP. The visual branch generates an initial prediction based on the visual features extracted by the image encoder. Subsequently, the cross-modal branch refines the initial prediction by addressing the discrepancy between the visual features and the textual semantics of the prediction. Since the image and text features are aligned in a joint image-text embedding space during pre-training, it becomes easy to identify this discrepancy. The cross-modal branch acts as a semantic-aware spell checker.
@@ III-C Decoder
The decoder aims to extract the character information from the visual feature $\bm{F}_{i}$ or cross-modal feature $\bm{F}_{c}$. The decoder framework is shown in Fig. 4. It adopts the design of the transformer decoder [23] plus the PSM technique mentioned in § III-A2, enabling a predicted character to have arbitrary dependencies on the input context during training.
The visual and cross-modal decoders have the same architecture but differ in the input. They receive the following inputs: a learnable position query $\bm{p}\in\mathbb{R}^{N\times D}$, an input context $\bm{c}\in\mathbb{R}^{N\times D}$, and a randomly generated attention mask $\mathcal{M}\in\mathbb{R}^{N\times N}$. $N$ represents the length of characters. The decoder outputs the prediction $\bm{y}\in\mathbb{R}^{N\times C}$, where $C$ is the number of character classes. The decoding stage can be denoted as: $\bm{y}=\texttt{DEC}(\bm{p},\bm{c},\mathcal{M},\bm{F})$. The first Multi-Head Attention (MHA) in Fig. 4 performs context-position attention:
@@
During training, the output of the decoder depends on the randomly permuted input context. This encourages the decoder to analyze the word structure beyond the traditional left-to-right or right-to-left sequence modeling assumptions [9]. The inclusion of a random attention mask $\mathcal{M}$ in Eq.(4) enables this capability [24]. Table I presents examples of generated attention masks, including a left-to-right auto-regressive (AR) mask, a cloze mask, and a random mask. Following PARSeq [24], we employ $K=6$ masks per input context during training. The first two masks are left-to-right and right-to-left masks, and others are randomly generated. CLIP4STR is optimized to minimize the sum of cross-entropy losses ($\texttt{CE}(\cdot)$) of the visual branch and the cross-modal branch:
@@
where $\hat{\bm{y}}$, $\bm{y}^{i}$, and $\bm{y}$ indicate ground truth, prediction of the visual branch, and prediction of the cross-modal branch.
@@ 22       $\bm{y}\leftarrow\texttt{Dec}^{c}(\bm{p}^{c},\bm{c},\mathcal{M}^{c},\bm{F}_{c})$;
Algorithm 1 Inference decoding scheme (§A)
@@ III-C1 Decoding scheme
CLIP4STR consists of two branches: a visual branch and a cross-modal branch. To fully exploit the capacity of both branches, we design a dual predict-and-refine decoding scheme for inference, inspired by previous STR methods [9, 24]. Alg. 1 illustrates the decoding process. The visual branch first performs autoregressive decoding, where the future output depends on previous predictions. Subsequently, the cross-modal branch addresses possible discrepancies between the visual feature and the text semantics of the visual prediction, aiming to improve recognition accuracy. This process is also autoregressive. Finally, the previous predictions are utilized as the input context for refining the output in a cloze-filling manner. The refinement process can be iterative. After iterative refinement, the output of the cross-modal branch serves as the final prediction.
IV Experiment
@@
IV-A Experimental Details
We instantiate CLIP4STR with CLIP-ViT-B/16, CLIP-ViT-L/14, and CLIP-ViT-H/14 [38]. Table II presents the main hyper-parameters of CLIP4STR. A reproduction of CLIP4STR is at https://github.com/VamosC/CLIP4STR.
Test benchmarks  The evaluation benchmarks include IIIT5K [74], CUTE80 [75], Street View Text (SVT) [76], SVT-Perspective (SVTP) [77], ICDAR 2013 (IC13) [21], ICDAR 2015 (IC15) [22], and three occluded datasets – HOST, WOST [78], and OCTT [14]. Additionally, we utilize 3 recent large benchmarks: COCO-Text (low-resolution, occluded text) [79], ArT (curved and rotated text) [15], and Uber-Text (vertical and rotated text) [16].
TABLE III: Word accuracy on 10 common benchmarks. Thebestandsecond-bestresults are highlighted. Benchmark datasets (B) - SVT, IIIT5K, IC13, and IC15. ‘N/A’ for not applicable. Reproduced by PARSeq [24].
Method Pre-train Data Train Data IIIT5K SVT IC13 IC15 IC15 SVTP CUTE HOST WOST OCTT
3,000 647 1,015 1,811 2,077 645 288 2,416 2,416 1,911